Please help

Express your answer using two significant figures.
Answer #1

M1 = 0.255 M H2SO4

V1 = 45 ml

n1 = M1*V1 = 0.255*(45/1000) = 0.01147

B)

M1 = 0.155 H2SO4

V1 = 190 ml

n1 = M1*V1 = 0.155*(190/1000) = 0.02945

C)

M1 = 0.120 H2SO4

V1 = 50 ml

n1 = M1*V1 = .120*50/1000 = 0.006